There are 9 black balls and 7 red balls in an urn. If 5 balls are drawn without replacement, what is the probability that exactly 3 black balls are drawn? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

sa Paga Here g black bellS 7 red kalls want to Bind the prahablity aure bl@( withaut replacemen Ya Remaining two ballc are ed
Answer #2
9C3*7C2/16C5= 84*21/4368=0.4038 I think
answered by: Yr